eviant:

The median age for a first marriage in the United States for men was 28.1 in 2009 and 28.2 in 2010. Use an exponential model to predict the median age for men in 2019, where x is the number of years since 2009.

eviant:

@Vocaloid

Vocaloid:

same logic as last time, you are given (0,28.1) and (1,28.2) y = ab^x, where a is the initial value 28.1 and b is the growth rate, x is the years since 2009 plugging this into y = ab^x gives us (28.2) = (28.1)b^(1), solve for b again, then re-write y = ab^x

eviant:

@Vocaloid 27.1?

Vocaloid:

if (28.2) = (28.1)b^1 what does b equal? you just have to divide both sides by 28.1

Vocaloid:

(28.2) = (28.1)b^1 b = 28.2/28.1

eviant:

27.0 27.1 29.3 29.2

eviant:

^^^those are the options

Vocaloid:

notice that the average age of marriage is increasing so it can't be 27.0 or 27.1 re-writing our exponential function gives us y = 28.1(28.2/28.1)^x 10 years passed between 2009 and 2019, so x = 10, plug in and solve for y

eviant:

29.2

Vocaloid:

yeah that's what i got too i'm not 100% sure but that's my best attempt